Solve the following problems. Show your solutions to support your answer.

1. At what rate will ₱ 6,500 accumulate to ₱ 27,000 for 3 years and 6 months, if compounded every 5 

months?

2. Accumulate ₱ 25,000 for 2 years and 5 months at 6 ½%, compounded every two months.

3. Briefly explain the difference between a “ 7% compounded monthly for a year” and a “7% simple 

interest for a year.”



1
Expert's answer
2021-04-14T14:11:56-0400

"A=P(1+r\/n)" nt

1.

"27000=6500(1+r\/2.4)"7.2

7.2 rt "(270\/65)=(1+r\/2.4)"

"r=0.524873784*100"

"r= 52.487%"


2.

"A=25000(1+0.065\/6)"29/6

"A=25000\times1.053459511"

"A=26336.48777"


3.

"Compound interest=Amount-Principal"

Assuming principal = 1

"A=P(1+r\/n)"nt

"A=1(1+0.07\/12)"12

"A=1.07229008"

Compound interest "= 1.07229008-1=0.07229008"


"Simple interest=P\times R \times T"

"=1 \times 0.07 \times 1=0.07"